# RMMI: Enhanced Obstacle Avoidance for Reactive Mobile Manipulation using an Implicit Neural Map

### v1

@@ V-B Results
The results of the simulation experiments are presented in Table I. Our approach, which integrates collisions constraints and a secondary cost function, achieved the highest performance, avoiding collisions in all attempts. When comparing our approach using variations on the robot representation by using different sets of sampled points alongside a robot representation using manually placed spheres. The approach reached a success rate of 88.5% for a point representation with 9476 points (with 11.5% getting stuck in local minima), followed by the manually placed spheres with 85.8%. When reducing the number of sampled points, the performance gradually reduces, with a 79.5% success rate when using 2358 points (see also Fig. 4).
@@ V-C1 Precision
Figure 4: Comparison of point representations using different levels of precision, where a faster control loop frequency is achieved at the cost of reduced performance in our benchmark tasks. The figure shows that as the number of sampled points decreases, the success rate drops, but the control loop operates at a higher frequency. Additionally, our approach using points instead of spheres results in smoother motions, as indicated by the lower average end-effector acceleration.
We evaluated our approach using different levels of precision in the robot representation, achieved by varying the total number of sampled points. As shown in Fig. 4, reducing the number of sampled points increased the controller frequency at the cost of performance. The reduced precision required a higher stopping distance to prevent interpenetration, resulting in less information being available to the control problem. For our real robot experiments, we selected the 2358 point representation, which offered the fastest controller frequency without reducing the success rate below 75%.
@@ V-C2 Effects of Active Collision
We tested various combinations of our collision avoidance components, including a sweep on $\lambda_{c}^{max}$ to control the gain of the active collision cost. These combinations were tested on an additional 100 variations of both simulated environments to avoid overlap between data used for hyperparameter tuning and method evaluation. The results are shown in Fig. 5, where the dashed lines represent the baseline without the active collision cost, the dotted lines show the performance with only active collision cost, and the solid lines depict our full proposed approach. The results indicate that using both the inequality constraints and the active collision cost improved the success rate across different sampled point variations and the sphere representation. However, $\lambda_{c}^{max}\geq 1.5$ led to diminishing returns, as the robot began prioritising obstacle avoidance over moving towards the goal.
@@
Figure 6: (Top) Reconstructed meshes of the environments used for the real-world experiments: (a) Table, (b) Bookshelf, and (c) Cabinet. The green markers indicate the target 6-DoF poses for the robot’s end effector in each scenario. (Bottom) Examples of the starting (blue) and final configurations of the robot during the experiments.
@@ VI Real Robot Experiments
VI-A Robot Representation
The robot representation used in the real-world experiments differs slightly from the simulation due to additional hardware considerations. Specifically, the real robot includes a RealSense D435 camera mounted on the end effector for capturing posed depth images, as well as a safety laser on the robot base. These additions required the inclusion of extra points and spheres to account for the camera and to extend the base representation, ensuring that the safety region where the mobile base operates is adequately covered.
VI-B Tasks and Setup
We evaluated our method in three real-world scenarios, as shown in Figure 6, with further details available in the supplementary video and on our website [^4] https://rmmi.github.io. Each scene was reconstructed using iSDF [23] to generate the neural SDF from 30 pairs of posed depth images captured by the robot.
The experiments involved a reaching task with pre-defined target end-effector poses in three scenes. For the Cabinet and Bookshelf scenes, we tested four target poses, while the Table scene involved eight target poses. Each target pose was tested five times, resulting in a total of 80 trials per experiment. In each trial, the robot started with the arm in a home configuration, and the base was positioned randomly in front of the scene.
@@ VI-C Results
Some examples of the starting configurations and target poses are shown in Fig. 6 and the attached video. The results across the three different scenes are presented in Table II. Our approach, which incorporates both the inequality constraints and active collision cost, achieved the highest performance, with 96.3% and 97.5% for the point-based and sphere-based representations, respectively. Importantly, our method did not result in any collisions, with failures occurring only due to local minima.
The addition of the active collision cost significantly improved performance by at least 12.5% for both point-based and sphere-based robot representations. Without collision avoidance, the robot collided in 45% of the runs, with an additional 25% of failures due to the robot becoming stuck when obstacles intruded the safety zone of the base.
Interestingly, the end-effector acceleration measurements revealed a 10% reduction in acceleration when using the sphere-based representation compared to the point-based one. This contrasts with the simulation results, where the opposite was the case. We expect that this is due to the higher refresh rate of the real-world controller with the sphere representation. In the simulations, the control loop was executed at a fixed rate across all representations.
@@ VII Conclusions
In this letter, we presented a reactive controller capable of reliably reaching target poses while avoiding static obstacles, achieving a success rate of 96% in real-world tasks. Leveraging a neural SDF, our approach efficiently integrates collision avoidance into the reactive controller, while maintaining a fixed control rate. In addition, we introduced a simple yet effective active collision cost, which is robust to parameter variations and increases the performance by at least 10% in both simulation and real-world experiments.
Our approach effectively generates collision-free motions in a holistic manner, demonstrating robustness and efficiency. Limitations of our approach include the shortsightedness inherent in the local reactive controller paradigm, which makes our approach susceptible to local minima. This limitation could be mitigated by complementing our controller with a global planning strategy that updates at a lower frequency. Moreover, the parallel computation of distances enables our controller to operate a receding horizon framework.
Our approach opens promising avenues for future research. One potential direction is to eliminate the assumption of a predefined camera path for reconstruction [23, 24]. Integrating our controller with advances in the the Neural Fields community [38] could allow for the autonomous generation of these representations using a robot, paving the way for more dynamic and adaptive control strategies.
In conclusion, our reactive controller offers a robust and efficient solution for avoiding collisions with static obstacles, while also providing a foundation for future research that could further enhance its capabilities and applicability in complex robotic tasks.

### v2

@@ V-B Results
The results of the simulation experiments are presented in Table I. Our approach, which integrates collisions constraints and a secondary cost function, achieved the highest performance, avoiding collisions in all attempts. When comparing different robot representations, our approach showed strong performance. Using a point representation with 9476 points, we achieved a success rate of 88.5% (with the remaining 11.5% getting stuck in local minima), followed by the manually placed spheres approach with 85.8%. When reducing the number of sampled points, the performance gradually reduces, with a 79.5% success rate when using only 2358 points (see also Fig. 4).
@@ V-C1 Precision
Figure 4: Comparison of point representations with varying number of sample points, where a faster control loop frequency is achieved at the cost of reduced performance in our benchmark tasks. The figure shows that as the number of sampled points decreases, the success rate drops, but the control loop operates at a higher frequency. Additionally, our approach using points instead of spheres results in smoother motions, as indicated by the lower average end-effector acceleration.
We evaluated our approach using different levels of precision in the robot representation, achieved by varying the total number of sampled points. As shown in Fig. 4, reducing the number of sampled points increased the controller frequency at the cost of performance. The reduced precision required a higher stopping distance to prevent interpenetration, resulting in less information being available to the optimisation problem. For our real robot experiments, we selected the 2358 point representation, which offered the fastest controller frequency without reducing the success rate below 75%.
@@ V-C2 Effects of Active Collision
We tested various combinations of our collision avoidance components, including a sweep on $\lambda_{c}^{\max}$ to control the gain of the active collision cost. These combinations were tested on an additional 100 variations of both simulated environments to avoid overlap between data used for hyperparameter tuning and method evaluation. The results are shown in Fig. 5, where the dashed lines represent the baseline without the active collision cost, the dotted lines show the performance with only active collision cost, and the solid lines depict our full proposed approach. The results indicate that using both the inequality constraints and the active collision cost improved the success rate across different sampled point variations and the sphere representation. However, $\lambda_{c}^{\max}\geq 1.5$ led to diminishing returns, as the robot began prioritising obstacle avoidance over moving towards the goal.
@@
Figure 6: (Top) Reconstructed meshes of the environments used for the real-world experiments: (a) Table, (b) Bookshelf, and (c) Cabinet. The green markers indicate the target 6-DoF end effector poses. (Bottom) Examples of the starting (blue) and final configurations of the robot during the experiments.
@@ VI Real Robot Experiments
VI-A Robot Representation
The robot representation used in the real-world experiments differs slightly from the simulation due to additional hardware considerations. These changes consider the geometry of the RealSense D435 mounted on the end effector and an inflated robot base to account for the Omron LD-60’s internal safety margins.
VI-B Tasks and Setup
We evaluated our method in three real-world scenarios, as shown in Figure 6, with further details available in the supplementary video and on our website: https://rmmi.github.io. Each scene was reconstructed using iSDF [23] to generate the neural SDF from 30 pairs of posed depth images captured by the robot.
During both the capture of the scene and the reaching task, the robot was localised in a 2D map using the in-built Omron-LD60 navigation capabilities.
The experiments involved a reaching task with pre-defined target end-effector poses in three scenes. For the Cabinet and Bookshelf scenes, we tested four target poses, while the Table scene involved eight target poses. Each target pose was tested five times, resulting in a total of 80 trials per experiment. In each trial, the robot started with the arm in a home configuration, and the base was positioned randomly in front of each scene.
@@ VI-C Results
Some examples of the starting configurations and target poses are shown in Fig. 6 and the attached video. The results across the three different scenes are presented in Table II. Our approach, which incorporates both the inequality constraints and active collision cost, achieved the highest performance, with 96.3% and 97.5% for the point-based and sphere-based representations, respectively. Importantly, our method did not result in any collisions, with the remaining failures occurring only due to local minima.
We observed the same trend regarding our active collision cost, observing an improvement in performance by at least 12.5% for both point-based and sphere-based robot representations. Without collision avoidance, the robot collided in 45% of the runs, with an additional 25% of failures due to the robot becoming stuck when obstacles entered the safety zone of the base.
Interestingly, the end-effector acceleration measurements revealed a 10% decrease in acceleration when using the sphere-based representation compared to the point-based one. This contrasts with the simulation results, which showed higher acceleration for the sphere-based representation. We suspect this discrepancy may be due to the higher refresh rate of the real-world controller when using the sphere representation, whereas in simulation, a fixed rate was used across all experiments.
@@ VII Conclusions
In this paper, we present a reactive controller capable of reliably reaching target poses while avoiding static obstacles, achieving a success rate of 96% in real-world tasks. Leveraging a neural SDF, our approach efficiently integrates collision avoidance into the reactive controller while maintaining a fixed control rate. In addition, we introduce a simple yet effective active collision cost, which is robust to parameter variations and increases performance by at least 10% in both simulation and real-world experiments.
Our approach effectively generates collision-free motions holistically, demonstrating robustness and efficiency. However, it is limited by the shortsightedness inherent in the local reactive controller paradigm, which makes it susceptible to local minima. This limitation could be mitigated by complementing our controller with a global planning strategy that updates at a lower frequency. Moreover, the parallel computation of distances enables our controller to operate in a receding horizon framework.
Our approach opens promising avenues for future research. One potential direction is to eliminate the assumption of a predefined camera path for reconstruction [23, 24]. Integrating our controller with advances in the Neural Fields community [38] could allow for the autonomous generation of these representations using a robot. Additionally, recent work has explored novel distance fields that can be efficiently updated in the presence of dynamic obstacles [39]. While the core principles of our approach remain applicable, evaluating its performance in such dynamic scenarios is left for future work.
In conclusion, our reactive controller offers a robust and efficient solution for avoiding collisions with static obstacles. The approach provides a foundation for future research that could further enhance its capabilities and applicability in complex robotic tasks.